OpenStudy (anonymous):

how would i begin to prove the equation 2/1+cosx - tan^2 x/2=1

OpenStudy (xapproachesinfinity):

so it is \[\frac{1}{1+\cot x}-\tan^2x=1\]

OpenStudy (anonymous):

okay i understand

OpenStudy (xapproachesinfinity):

the first thing im thinking about is multiplying by 1-cotx top and bottom for the 1/1+cotx

OpenStudy (xapproachesinfinity):

i haven't started yet i was just asking is that the right thing? because the way you wrote it can be misleading

OpenStudy (xapproachesinfinity):

oh hold on i missed 2 on top

OpenStudy (anonymous):

well it's like \[(2/1+cosx) - \tan^2 x/2=1\]

OpenStudy (xapproachesinfinity):

\[\frac{2}{1+\cot x}-\frac{\tan^2x}{2}=1\]

OpenStudy (xapproachesinfinity):

may bad! cos not cot

OpenStudy (anonymous):

right right i got it

OpenStudy (xapproachesinfinity):

i must be blind or something lol

OpenStudy (xapproachesinfinity):

\[\frac{2}{1+\cos x}-\frac{\tan^2x}{2}=1\] looking at the left side \[\frac{2}{1+\cos x}-\frac{\tan^2x}{2}=\frac{2(1-\cos x)}{(1+\cos x)(1-\cos x)}-\frac{\tan^2x}{2}\]

OpenStudy (xapproachesinfinity):

let's go from there

OpenStudy (xapproachesinfinity):

\[\frac{2(1-\cos x)}{1-\cos^2x}-\frac{\tan^2x}{2}\] \[\frac{2-2\cos x}{\sin^2x}-\frac{\tan^2x}{2}\] \[\frac{2-2\cos x}{\sin^2x}-\frac{\sin^2x}{2\cos^2x}\]

OpenStudy (xapproachesinfinity):

can you go further your self

OpenStudy (xapproachesinfinity):

actually that's not an identity

OpenStudy (anonymous):

i didnt think so lol

OpenStudy (xapproachesinfinity):

oh no i take it back it is! when i plugged in 0 i forgot something lol

OpenStudy (xapproachesinfinity):

hmm for pi/4 it is not good!

OpenStudy (xapproachesinfinity):

are you sure you typed the right thing!

OpenStudy (anonymous):

i will show you exactly what's in the textbook but how do you get the horizontal division line in the equation like you did in all your formulas?

OpenStudy (xapproachesinfinity):

use drawing tool if you have a mouse you will do find drawing the problem

OpenStudy (xapproachesinfinity):

snap shot your problem don't you have a phone?

OpenStudy (xapproachesinfinity):

what you have there is not an identity i verified it with couple values and it fails miserably if it is an identity it should not feel even one

OpenStudy (xapproachesinfinity):

fail*

OpenStudy (anonymous):

\[\frac{ 2 }{ 1 + cosx } - \tan^2\frac{ x }{ 2 } = 1\]

OpenStudy (xapproachesinfinity):

ah wow why you didn't correct me above i was dividing tan^2 over 2 2 was supposed to be within tan

OpenStudy (xapproachesinfinity):

darn you made me work really hard lol

OpenStudy (anonymous):

im sorry i wouldve if i known it was wrong but i was confused on this entire problem hence why i am here asking for help

OpenStudy (xapproachesinfinity):

here what you can do \[\frac{2}{1+\cos x}=1+\tan^2(x/2)=\sec^2(x/2)\] if you prove the left is sec^2(x/2) then you are done!

OpenStudy (xapproachesinfinity):

that is the same as proving \[\frac{1+\cos x}{2}=\cos^2(x/2) \]

OpenStudy (xapproachesinfinity):

isn't that some familiar identity to yoou?

OpenStudy (xapproachesinfinity):

the last line a i wrote should be obvious to you

OpenStudy (anonymous):

i get it i think the sec and tan identity was what i was thinking i would get to to prove the identity i just wasnt sure how to get there

OpenStudy (xapproachesinfinity):

1+cosx/2=cos^2(x/2) is called half angle identity which can be proven easily but you don't need to draw it all the way as long as that is an identity we know of so the initial problem is an identity as well

OpenStudy (anonymous):

right right i remember that now

OpenStudy (xapproachesinfinity):

let's prove that half angle identity for the sake of argument we know that cosx=2cos^2(x/2)-1 so 1+cosx=2cos^2(x/2) cos^2(x/2)=(1+cosx)/2 hence proved [cos(2a)=2cos^2a-1 is another useful identity]

OpenStudy (xapproachesinfinity):

they are all linked!

OpenStudy (xapproachesinfinity):

oh well gotta leave!

OpenStudy (anonymous):

how do i prove it's sec though?

OpenStudy (anonymous):

@zepdrix I was wondering if you can demonstrate this identity 2/1+cosx - tan^2 x/2=1

OpenStudy (anonymous):

I'm pretty new to this myself, so I'm trying to understand what @xapproachesinfinity did but I think I'm kinda lost...

zepdrix (zepdrix):

Mmm can we do it in a new thread? :D I don't feel like reading through all of these comments XD lol

OpenStudy (anonymous):

mm I pretty much a noob in trig identity LOL but ok xD

OpenStudy (xapproachesinfinity):

@Learner1298 we already proved it? why you need to prove sec lol read through what i did

OpenStudy (loser66):

\[\dfrac{2}{1+cos (x) }-tan^2(x/2)\] \[= \dfrac{2}{1+cos(x)} -\dfrac{sin^2(x)}{(1+cos (x))^2}\] because \(tan (x/2) = \dfrac{sin (x) }{1+cos x}\)

OpenStudy (loser66):

hence, it turns to \(\dfrac{2(1+cos(x)-sin^2(x)}{(1+cos(x))^2}\) Now, numerator only:

OpenStudy (loser66):

\(2(1+ cos(x) ) -sin^2(x)\\=2+2cosx-(1-cos^2(x) =2+2cos(x)-1+cos^2(x) \\=1+2cos(x) +cos^2(x) = (1+cos(x))^2 \)

OpenStudy (loser66):

Put it back to the fraction above, we can see that numerator = denorminator , hence the fraction =1 = R H S

OpenStudy (loser66):

@xapproach.... hey kid, the second reader didn't understand your stuff. Let old man helps you out. hahahaha.....................

OpenStudy (xapproachesinfinity):

you did a different version lol this trig stuff there are often many diff ways of doing stuff